### v1

# Boolean-Narayana numbers

Abstract
We introduce a refinement of Boolean-Catalan numbers and call them Boolean-Narayana numbers. We provide an explicit formula for these numbers, and prove unimodality, log-concavity, and real-roots-only results for their sequences.
1 Introduction
A binary plane tree is a rooted plane tree on unlabeled vertices in which every non-leaf vertex has one or two children, and every child is a left child or a right child, even if it is the only child of its parent. It is well-known that the number of such trees on $n$ vertices is the $n$th Catalan number $C_{n}={2n\choose n}/(n+1)$. We will consider an enhanced version of these trees.
Definition 1.1
A 0-1-tree is a binary plane tree in which every vertex that has two children is labeled 0 or 1.
Figure 1 shows the six 0-1-trees on three vertices.
Figure 1: The six 0-1 trees on three vertices.
The number of 0-1-trees on $n$ vertices is denoted by $B(n)$ and is called a Boolean-Catalan number [10]. They count various objects, some of which are discussed in the thesis of Chetak Hossain [10], and some others appear in papers of Colin Defant [7, 8, 9], later to be shown to bijectively correspond to 0-1-trees by the present author [3]. The name ”Boolean-Catalan numbers” can be justified, for example, by saying that binary plane trees are counted by the Catalan numbers, and if a binary plane tree has $m$ vertices with two children, then we associate an element of the Boolean algebra $\{0,1\}^{m}$ to it.
In this paper, we take this concept further. The Narayana numbers are the best-known refinement of the Catalan numbers. Accordingly, they have many equivalent definitions. The one that is closest to our subject defines the Narayana number $N(n,k)$ as the number of binary plane trees on $n$ vertices with k1k-1 right edges. So we define the Boolean-Narayana numbers $BoNa(n,k)$ as the number of 0-1-trees on $n$ vertices that have k1k-1 right edges.
In Section 2, we provide two additional combinatorial interpretations for the Boolean-Narayana numbers. In Section 3, we use the Lagrange Inversion Formula to prove an explicit formula for these numbers. Then in Section 4, we construct an injection to prove that the for fixed $n$, the sequence of the numbers $BoNa(n,k)$, where $k=1,2,\cdots,n$, is unimodal. We follow this by proving in Section 5 that the generating polynomial of that sequence has real roots only, which implies that it is log-concave. In Section 6, we fix $k$, and prove that the infinite sequence $BoNa(n,k)$, where $n=k+1,k+2,\cdots$, is log-concave. That is, we will see that the Boolean-Narayana numbers share the best-known properties of the Narayana numbers. We end by discussing future directions in Section 7, which also explain our original motivation in this family of questions.
2 A combinatorial interpretation
Stack sorting or West stack sorting of permutations is an interesting and well-studied field of research. See Chapter 8 of [2] for an introduction. The stack sorting operation $s$ has three natural, equivalent definitions. Here we mention one of them. We define $s$ on all finite permutations $p$ and their substrings by setting $s(1)=1$ for the one-element permutation, and setting $s(p)=s(LnR)=s(L)s(R)n$, where $n$ is the maximum entry of the string $p$, while $L$ is the substring of entries on the left of $n$, and $R$ is the substring on the entries on the right of $n$. Note that $L$ and $R$ can be empty.
The questions, methods and results about the stack sorting algorithm are intrinsically connected to the theory of pattern-avoiding permutations, a large and rapidly evolving field. We say that a permutation $p$ contains the pattern (or subsequence) $q=q_{1}q_{2}\cdots q_{k}$ if there is a $k$-element set of indices $i_{1}<i_{2}<\cdots<i_{k}$ such that $p_{i_{r}}<p_{i_{s}}$ if and only if $q_{r}<q_{s}$. If $p$ does not contain $q$, then we say that $p$ avoids $q$. For example, $p=3752416$ contains $q=2413$, as the first, second, fourth, and seventh entries of $p$ form the subsequence 3726, which is order-isomorphic to $q=2413$. A recent survey on permutation patterns by Vatter can be found in [12].
Recently, in a series of papers [7, 8, 9] Colin Defant studied the size of $s^{-1}(Class_{n})$, that is, he studied the number of permutations $p$ of length $n$ for which $s(p)$ belongs to a certain permutation class. Let $\operatorname{Av}_{n}(q)$ denote the set of permutations of length $n$ that avoid the pattern $q$. The classic result that the number of permutations $p$ of length $n$ for which $s(p)$ is the identity permutation is the Catalan number $C_{n}$ is a special case of this line of work. Namely, it is the result that $|s^{-1}(\operatorname{Av}_{n}(21))|=C_{n}$. Considering various permutation classes defined by sets of patterns of length three, Defant has proved the chain of equalities
$\sum_{n\geq 1}|s^{-1}(\operatorname{Av}_{n}(132,312))|z^{n}$ $=$ $\sum_{n\geq 1}|s^{-1}(\operatorname{Av}_{n}(231,312))|z^{n}$
$=\sum_{n\geq 1}|s^{-1}(\operatorname{Av}_{n}(132,231))|z^{n}$ $=$ $\frac{1-2z-\sqrt{1-4z-4z^{2}}}{4z}.$
In [3], the present author showed the equality $\sum_{n\geq 1}B(n)z^{n}=\frac{1-2z-\sqrt{1-4z-4z^{2}}}{4z}$, and then he constructed a bijection (of a recursive nature) from the set $s^{-1}(\operatorname{Av}_{n}(231,312))$ of permutations to the set $\mathcal{B}_{n}$ of 0-1-trees on $n$ vertices, and another bijection from the set $s^{-1}(\operatorname{Av}_{n}(132,312))$ of permutations to $\mathcal{B}_{n}$. In this section, we will use the ideas presented there to establish combinatorial interpretations of the Boolean-Narayana numbers. If $p=p_{1}p_{2}\cdots p_{n}$ is a permutation, then we say that $i$ is an ascent of $p$ if $p_{i}<p_{i+1}$. If $p_{i}>p_{i+1}$, then we say that $i$ is a descent of $p$.
Proposition 2.1
The number of permutations in $s^{-1}(\operatorname{Av}_{n}(231,312))$ that have exactly k1k-1 descents is $BoNa(n,k)$.
Proof: Let $p=LnR$ be a permutation of length $n$. For $s(p)$ to avoid both 231 and 312, it is necessary that both $L$ and $R$ avoid those two patterns. If $L$ and $R$ are both nonempty, then it is proved in [3] that regardless of what $L$ and $R$ are, there are exactly two possible choices for the underlying set of $L$ (the underlying set of $R$ is determined by these choices) so that $s(p)\in\operatorname{Av}_{n}(231,312)$. If $L$ (resp. $R$) is empty, then the sufficient and necessary condition for $s(p)\in\operatorname{Av}_{n}(231,312)$ is that $s(R)$ (resp $s(L)$) avoids both 231 and 312.
Let $A(n,k)$ be the number of permutations in $s^{-1}(\operatorname{Av}_{n}(231,312))$ that have exactly k1k-1 descents. Then $A(1,1)=1$, and it follows from the analysis in the previous paragraph that
$A(n,k)=A(n-1,k)+A(n-1,k-1)+2\sum_{i=2}^{n-1}\sum_{j\geq 0}A_{i-1,j}A_{n-i,k-1-j}.$ (1)
Indeed, on the right-hand side, the first summand counts permutations in which $n$ is in the last position, the second summand counts those in which $n$ is in the first position, and the third summand counts those where $n$ is in the middle and counts them according to the number j1j-1 of descents in $L$.
On the other hand, $B(1,1)=1$, the numbers $BoNa(n,k)$ obviously satisfy the analogous recurrence relation
$BoNa(n,k)=BoNa(n-1,k)+BoNa(n-1,k-1)+2\sum_{i=2}^{n-1}\sum_{j\geq 0}BoNa(i-1,j)BoNa(n-i,k-1-j),$ (2)
since the number of right edges in a 0-1 tree is easy to count by considering subtrees of the root. This proves our claim. $\Diamond$
Proposition 2.2
The number of permutations in $s^{-1}(\operatorname{Av}_{n}(132,312))$ that have exactly k1k-1 descents is $BoNa(n,k)$.
Proof: Analogous to the proof of Proposition 2.1, with the needed statement on the underlying sets of $L$ and $R$ proved in [3]. $\Diamond$
Note that it follows by reflection through a vertical axis that BoNa(n,k)=BoNa(n,nk+1)BoNa(n,k)=BoNa(n,n-k+1). So Propositions 2.1 and 2.2 yield the following.
Corollary 2.3
For all $n$ and $k$, the equality A(n,k)=A(n,nk+1)A(n,k)=A(n,n-k+1) holds. That is, the number of permutations in $s^{-1}(\operatorname{Av}_{n}(231,312))$ with k1k-1 ascents is equal to the number of permutations in $s^{-1}(\operatorname{Av}_{n}(231,312))$ with k1k-1 descents. The same holds for $s^{-1}(\operatorname{Av}_{n}(132,312))$.
3 Enumeration
In this section, we will use the Lagrange Inversion Formula to find an explicit formula for $BoNa(n,k)$. The Lagrange Inversion Formula, in its simplest form, can be found in many graduate-level textbooks. However, in this section, we will use its univariate version on a function that has two variables, and for that reason, we will explain which version we use and how.
The Lagrange Inversion Formula, in one of its univariate versions, is the following.
Theorem 3.1
Let $F(z)$ be the unique formal power series with $F(0)=0$ satisfying
$F(z)=z\phi(F(z),$ (3)
where $\phi(y)$ is a formal power series with $\phi(0)\neq 0$. Then for all positive integers $n$, the equality
$z^{n}]F(z)=\frac{1}{n}[y^{n-1}\phi(y)^{n}]$
holds.
Note that the name ”inversion formula” is justified because (3) is equivalent to the fact that $F$, as a function of $z$, is the compositional inverse of the function $z/\phi(z)$.
Now let us assume that we have a bivariate power series $F(z,u)$ that is defined implicitly by the equation
$F(z,u)=z\phi(F(z,u),u),$ (4)
where $\phi(0,u)\neq 0$. Then we can treat $u$ as a parameter and apply the univariate formula (3) in the variable $z$, since (5) is equivalent to stating that the compositional inverse of $F(z,u)$, in the variable $z$, is equal to $z/\phi(F(z,u),u)$.
As BoNa(n,k)=BoNa(n,nk+1)BoNa(n,k)=BoNa(n,n-k+1), it suffices to compute $BoNa(n,k)$ for the case when $k\leq(n+1)/2$.
Theorem 3.2
Let $k\leq(n+1)/2$, where $n$ and $k$ are positive integers. Then
$BoNa(n,k)=\frac{1}{n}{n\choose k-1}\sum_{j=0}^{k-1}2^{j}{k-1\choose j}{n-k+1\choose j+1}$.
Proof: Let $t(n,k)=BoNa(n,k+1)$ be the number of 0-1 trees on $n$ vertices with $k$ right edges, and let
$T(z,u)=\sum_{n\geq 1,k\geq 0}t(n,k)z^{n}u^{k}=1+\sum_{n\geq 1,k\geq 1}BoNa(n,k)z^{n}u^{k-1}.$ (5)
Removing the root of a 0-1 tree we either get the empty set, or one tree that was connected to the root by a left edge, or one tree that was connected to the root by a right edge, or two trees, one of which was connected to the root by a left edge. This implies the functional equation
$T=z(1+uT+T+2uT^{2}).$
In order to use the Lagrange Inversion Formula, we set
$\phi(y)=\phi(y,u)=1+(u+1)y+2uy^{2},$ (6)
and then (6) becomes
$T=z\phi(T)).$ (7)
By Theorem 3.1,
$z^{n}u^{k-1}]T(z,u)=\frac{1}{n}[y^{n-1}u^{k-1}]\phi(y)^{n}.$ (8)
Now notice that
$\phi(y)=(1+y)+u(y+2y^{2})=(1+y)+uy(1+2y).$
Therefore, for $\phi(y)^{n}$ on the right-hand side of (8), we take the power $((1+y)+uy(1+2y))^{n}$, and notice that the term $u^{k-1}$ will come from the summand of that binomial expansion in which the $u$-term is taken to the (k1)(k-1)st power. That is,
$u^{k-1}]\phi(y)^{n}={n\choose k-1}(1+y)^{n-k+1}y(1+2y))^{k-1}.$
Recalling (8), this means
$BoNa(n,k)=[z^{n}u^{k-1}]T(z,u)=\frac{1}{n}{n\choose k-1}[y^{n-1}]y^{k-1}\cdot(1+2y)^{k-1}(1+y)^{n-k+1},$
or, equivalently,
$BoNa(n,k)=[z^{n}u^{k-1}]T(z,u)=\frac{1}{n}{n\choose k-1}[y^{n-k}](1+2y)^{k-1}(1+y)^{n-k+1}.$
Finally, expanding the binomial sums we get the expression
$BoNa(n,k)=\frac{1}{n}{n\choose k-1}\sum_{j=0}^{\min(k-1,n-k)}2^{j}{k-1\choose j}{n-k+1\choose j+1}$,
However, recall that we assumed that $k\leq(n+1)/2$, (which, for symmetry reasons, does not result in loss of generality), so $k-1\leq n-k$, and our theorem is proved. $\Diamond$
The numbers $BoNa(n,k)$ for $1\leq n\leq 5$, and $k=0,1,\cdots,n-1$ are shown below. Note the interesting fact that $BoNa(n,2)=(n-1)^{2}$.
  • 1
  • 1, 1
  • 1, 4, 1
  • 1, 9, 9, 1
  • 1, 16, 38, 16, 1.
4 Unimodality for fixed $n$
We say that sequence $a_{1},\cdots,a_{n}$ is unimodal if there exists an index $k$ so that $a_{1}\leq a_{2}\leq\cdots a_{k}\geq a_{k+1}\geq\cdots\geq a_{n}$. In this section we present a combinatorial proof of the following.
Theorem 4.1
For all fixed $n$, the sequence $BoNa(n,k)_{1\leq k\leq n}$ is unimodal.
Proof: As for any fixed $n$, the sequence $BoNa(n,k)_{1\leq k\leq n}$ is symmetric, it suffices to show that $BoNa(n,k)\leq B(n,k+1)$ if $k\leq(n-1)/2$. Let $\mathcal{B}oNa(n,k)$ be the set of 0-1 trees with k1k-1 right edges. We will construct an injection $z:\mathcal{B}oNa(n,k)\rightarrow\mathcal{B}oNa(n,k+1)$. Note that our injection $z$ is a slightly modified version that the present author used to prove a unimodality result on $t$-stack sortable permutations in [4, 5]. A perhaps more accessible source is Section 8.2.3 of [2].
In order to define this injection $z$, we need two tools. First, we define an involution $f:\mathcal{B}(n)\rightarrow\mathcal{B}(n)$, where $\mathcal{B}(n)$ is the set of all 0-1 trees on $n$ vertices, as follows. Let $T\in\mathcal{B}(n)$, and let $v$ be a vertex of $T$. Construct $f(T)$ as follows.
  1. (a)
    If $v$ has zero or two children, leave the subtrees of $v$ unchanged.
  2. (b)
    If $v$ has a left subtree only, then turn that subtree into a right subtree.
  3. (c)
    If $v$ has a right subtree only, then turn that subtree into a left subtree.
Leave all the labels unchanged. Let $f(T)$ be the tree we obtain from $T$ in this way. See Figure 2 for an example.
Figure 2: A 0-1 tree $T$ and its image $f(T)$.
Note that if $T$ has k1k-1 right edges, then $f(T)$ has k1k-1 left edges (and so, nkn-k right edges).
Second, for each $T\in\mathcal{B}(n)$, we define a total order of the nodes of $T$ as follows. Let us say that a node $v$ of $T$ is on level $j$ of $T$ if the distance of $v$ from the root of $T$ is $j$. Then our total order consists of listing the nodes on the highest level of $T$ going from left to right, then the nodes on the second highest level left to right, and so on, ending with the root of $T$.
Let $T_{i}$ be the subgraph of $T$ induced by the smallest $i$ vertices in this total order. Then $T_{i}$ is either a tree or a forest with at least two components. If $T_{i}$ is a tree, then $f(T_{i})$ is a tree as given by the definition of the map $f:\mathcal{B}(n)\rightarrow\mathcal{B}(n)$. If $T_{i}$ is a forest, then we define $f(T_{i})$ as the plane forest whose $h$th component is the image of the $h$th component of $T_{i}$.
Now let $k\leq(n-1)/2$, and let $T\in\mathcal{B}oNa(n,k)$. We define $z(T)$ as follows. Take the sequence $T_{1},T_{2},\cdots,T_{n}.$ Denote by $\ell(T)$ the number of left edges of the forest $T$, and by $r(T)$ the number of right edges of the forest $T$. Find the smallest index $i$ so that $\ell(T_{i})-r(T_{i})=1$. We will now explain why such an index always exists. If $T_{2}$ is a left edge, then $\ell(T_{2})-r(T_{2})=1-0=1$, and we are done. Otherwise, at the beginning we have $\ell(T_{2})-r(T_{2})<1$, while at the end we have
$\ell(T_{n})-r(T_{n})=(n-1-(k-1))-(k-1)=n-2k+1\geq 1,$
because of the restriction $k\leq(n-1)/2$. So at the beginning, $\ell(T_{j})-r(T_{j})$ is too small, while at the end, it is too large. On the other hand, it is obvious that as $i$ changes from 2 to $n$, at no step could $\ell(T_{j})-r(T_{j})$ “skip” a value as it could only change by 1. Therefore, for continuity reasons, it has to be equal to 1 at some point, and we set $i$ to be the smallest index for which this happens.
Now apply $f$ to the forest $T_{i}$, and leave the rest of $T$ unchanged. Let $z(T)$ be the obtained tree. Before we prove that $z$ is an injection, let us consider an example.
Example 4.2
Figure 3 shows an example of the injection $z:\mathcal{B}(9,4)\rightarrow\mathcal{B}(9,5)$. The large numbers denote the label of vertices with two children, whereas the small labels show the place of a vertex in the total order we defined above. In this example, the smallest index $i$ so that $\ell(T_{i})-r(T_{i})=1$ is $i=3$.
Figure 3: The action of $z$ on a tree in $\mathcal{B}(9,4)$.
Lemma 4.3
For all $k\leq(n-1)/2$, the function $z$ described above is an injection from $\mathcal{B}(n,k)$ into $\mathcal{B}(n,k+1)$.
Proof: It is clear that $z$ maps into $\mathcal{B}(n,k+1)$ as $z$ consists of the application of $f$ to a subgraph $T_{i}$ in which left edges outnumber right edges by one. We know that $f$ turns the number of left edges into the number of right edges and vice versa, so $z$ indeed increases the number of right edges by 1.
To see that $z$ is an injection, note that $z(T_{i})$ is the smallest subforest of $z(T)$ that consists of the first few nodes of $z(T)$ in the total order of all nodes and in which right edges outnumber left edges by one.
Now let $U\in\mathcal{B}(n,k+1)$. If $U$ does not have a subforest that consists of the first few nodes of $U$ in the total order of all nodes and in which the right edges outnumber the left edges by one, then by the previous paragraph, $U$ has no preimage under $z$. Otherwise, the unique preimage $z^{-1}(U)$ can be found by finding the smallest such subforest, applying $f$ to it, and leaving the rest of $U$ unchanged. $\Diamond$
Lemma 4.3 shows that $BoNa(n,k)\leq BoNa(n,k+1)$ if $k\leq(n-1)/2$. This implies that the sequence $\{BoNa(n,k)\}_{1\leq k\leq n}$ weakly increases to its midpoint. As the sequence is symmetric, it follows that it weakly decreases after that. Therefore, the sequence is unimodal as claimed. $\Diamond$
Note that $z$ never changes the labels of any vertex, nor does it change the ability of a vertex to receive a label, so this injection can be extended to other classes of labeled trees.
5 The real zeros property
Let $BoNa_{n}(u)=\sum_{k=1}^{n}BoNa(n,k)u^{k}$. We call these polynomials the Boolean-Narayana polynomials.
Theorem 5.1
For all positive integers $n$, the Boolean-Narayana polynomial $BoNa_{n}(u)$ has real roots only.
Proof: Note that $BoNa_{n}(u)=[z^{n}](uT(z,u))$. Note that for a fixed $u$, (8) implies, using (6), that
$z^{n}]T(z)=\frac{1}{n}[y^{n-1}]\phi(y)^{n},$
while if we view $u$ as a variable, then we get
$\sum_{k\geq 1}^{n}BoNa(n,k)u^{k-1}=[z^{n}]T(z,u)=\frac{1}{n}[y^{n-1}]\left(1+(1+u)y+2uy^{2}\right)^{n}.$
Multiplying both sides by $u$, we get
$BoNa_{n}(x)=\frac{u}{n}[y^{n-1}]\left(1+(1+u)y+2uy^{2}\right)^{n}.$ (9)
Now note that $\phi(y,u)$ is a quadratic polynomial in $y$ as we saw in (6). Let
$\phi(y,u)=(1+(1+u)y+2uy^{2}=(1-r_{1}y)(1-r_{2}y),$
so that $r_{1}+r_{2}=-(1+u)$ and $r_{1}r_{2}=2u$. Note that the discriminant of $\phi(y,u)$ as a quadratic polynomial of $y$ is $(1+u)^{2}-8u=1-6u+u^{2}$, which is positive if $u<0$, so $r_{1}$ and $r_{2}$ are real for negative $u$. Then (9) becomes
$BoNa_{n}(u)=\frac{u}{n}[y^{n-1}](1-r_{1}y)^{n}(1-r_{2}y)^{n}.$
Now let us expand and extract the coefficient of $y^{n-1}$ in the last displayed equation to get
$y^{n-1}](1-r_{1}y)^{n}(1-r_{2}y)^{n}=(-1)^{n-1}r_{2}^{n-1}\sum_{j=0}^{n-1}{n\choose j}{n\choose j+1}\left(\frac{r_{1}}{r_{2}}\right)^{j}.$ (10)
Recall that $r_{1}$ and $r_{2}$ are functions of $u$, and introduce the function $q(u)=r_{1}(u)/r_{2}(u)$. Also recall that for negative real $u$, both $r_{1}(u)$ and $r_{2}(u)$ are real numbers; therefore, so is $q(u)$.
The numbers ${n\choose j}{n\choose j+1}$ are constant multiples of the Narayana numbers $N(n,j)$. Therefore, we define the Narayana polynomials as
$N_{n}(q)=\frac{1}{n}\sum_{j=0}^{n-1}{n\choose j}{n\choose j+1}q^{j}.$ (11)
Using this notation, formula (10) turns into
$BoNa_{n}(u)=u(-1)^{n-1}r_{2}^{n-1}N_{n}(q(u)).$
It was proved by Francesco Brenti (see Theorem 5.3.1 in [6]) that all n1n-1 roots of $N_{n}(u)$ are real and negative. It is also known that the roots are simple (see Theorem 3.2 in [11]). The last displayed equation shows that if $q(u)$ is a root of $N_{n}$, then $u$ is a root of $BoNa_{n}$. As both polynomials have degree $n$ and have a root equal to 0, this will complete the proof of our theorem if we can show that the function $q(u)$ is bijective from the set $(-\infty,0)$ to itself. In order to do that we compute
$q(u)=\frac{r_{1}(u)}{r_{2}(u)}=\frac{1+u-\Delta}{1+u+\Delta},$
where $\Delta=\sqrt{1-6u+u^{2}}$. Then
$q^{\prime}(u)=\frac{-8(u-1)}{(u+1+\Delta)^{2}\Delta}.$
Now note that if $u<0$, then the 8(u1)=8(1u)>0-8(u-1)=8(1-u)>0, and $(u+1+\Delta)^{2}>0$, so $q^{\prime}(u)>0$, and so $q(u)$ is an increasing function on the interval $(-\infty,0)$. Finally, $q(u)$ takes all negative real values, since it is a continuous function on the set of negative reals and $\lim_{u\rightarrow-\infty}q(u)=$-$\infty$ and $\lim_{u\rightarrow 0}q(u)=0$. $\Diamond$
We say that sequence $a_{1},a_{2},\cdots$ of nonnegative real numbers is log-concave is log-concave if for all $i$, the inequality $a_{i-1}a_{k+1}\leq a_{i}^{2}$ holds. For finite sequences $a_{1},a_{2},\cdots,a_{n}$, it is well known [6] that if the generating polynomial $P(z)=\sum_{i=1}^{n}a_{i}z^{i}$ has real roots only, then the sequence $a_{1},a_{2},\cdots,a_{n}$ is log-concave. So Theorem 5.1 proves the following.
Corollary 5.2
For all fixed $n$, the sequence $BoNa(n,1),BoNa(n,2),\cdots,BoNa(n,n)$ is log-concave.
In other words, the Boolean-Narayana numbers are ”horizontally log-concave”. It would be interesting to give a combinatorial proof of this fact. We note that finite log-concave sequences of positive real numbers are unimodal, so it is perhaps not surprising that a combinatorial proof of Corollary 5.2 is more difficult to find than such a proof of Theorem 4.1.
6 Log-concavity for fixed $k$
Theorem 6.1
For any fixed $k>1$, the sequence $\{BoNa(n,k)\}_{n\geq k+1}$ is log-concave.
In other words, the Boolean-Narayana numbers are ”vertically log-concave”. We will need the following lemma, which is Proposition 2.5.1 in [6]. Note that we say that a sequence of real numbers $z_{0},z_{1},\cdots$ has no internal zeros if there are no three indices $i<j<k$ so that $z_{i}z_{k}\neq 0$ but $z_{j}=0$.
Lemma 6.2
Let $a$ and $b$ be two nonnegative integers, and let $z_{0},z_{1},\cdots$ be a log-concave sequence of nonnegative real numbers with no internal zeros. Then the sequence $v_{0},v_{1},\cdots$ defined by
$v_{n}=\sum_{j=0}^{n}{n\choose j}z_{j}=\sum_{j\geq 0}{n\choose j}z_{j}$
is log-concave.
This result is often mentioned by saying that the binomial transform preserves log-concavity as the expression on the right-hand side is the binomial transform of the sequence $z_{n}$.
In order to be able to use Lemma 6.2, we rearrange the expression obtained for $BoNa(n,k)$ in Theorem 3.2 so that it represents the binomial transform of a log-concave sequence. We achieve this as follows. First, we note that $\frac{1}{n}{n\choose k-1}=\frac{1}{k-1}{n-1\choose k-2}$. This turns the formula from Theorem 3.2 into
$BoNa(n,k)=\frac{1}{k-1}{n-1\choose k-2}\sum_{j=0}^{k-1}2^{j}{k-1\choose j}{n-k+1\choose j+1}$. (12)
Note that after expanding the product above, we can make good use of the identity
${n-1\choose k-2}{n-k+1\choose j+1}={k+j-1\choose k-2}{n-1\choose k+j-1}$.
Indeed, it is routine to check that the two sides are identical by simply using the definition of binomial coefficients. The advantage of the format on the right-hand side is that only one of the factors there depends on $n$. Comparing the last displayed equation with (12), we see that
$BoNa(n,k)=\sum_{j=0}^{k-1}a_{j}{n-1\choose k+j-1}$, (13)
where
$a_{j}=\frac{2^{j}}{k-1}{k-1\choose j}{k+j-1\choose k-2}$.
Note that the $a_{j}$ depend on $k$ and $j$ only, and not on $n$.
Let us define a sequence $c_{m}$ so that $c_{m}=a_{m-(k-1)}$ if $k-1\leq m\leq 2k-2$, and $c_{m}=0$ otherwise. Then setting j=m(k1)j=m-(k-1), we see that (13) becomes
$BoNa(n,k)=\sum_{m\geq 0}c_{m}{n-1\choose m}$,
and our proof is complete by Lemma 6.2 with n1n-1 playing the role of $n$.
7 Further directions
It is well-known that unlabeled binary plane trees are in natural bijection with 231-avoiding permutations of length $n$. On the other hand, decreasing binary trees, that is, plane binary trees whose vertices are bijectively labeled with the number $\{1,2,\cdots,n\}$ so that each vertex has a smaller label than its parent, are in natural bijection with all permutations of length $n$. Therefore, these trees are counted, respectively, by the Catalan numbers and factorials. In both cases, we can refine our counting argument to focus on trees with a given number of right edges, and then we get, respectively, the Narayana numbers and the Eulerian numbers.
In this paper, we have shown that 0-1 trees are an intermediate class of trees between those two mentioned in a previous paragraph, and we provided combinatorial interpretations for them in terms of permutations. We hope to extend on this in a future paper by finding other natural classes of trees which are in natural bijections with permutations so that the number of right edges of the trees corresponds to the number of descents of the permutations.